OpenStudy (anonymous):

(sin^-1){(x^2)/[(sqrt(x^4)+(a^4))]} differentiate w.r.t. x

OpenStudy (anonymous):

we can substitute \[x ^{2} = a ^{2}\sin \theta\]

OpenStudy (mimi_x3):

\[\sin^{-1}\left( \frac{x^{2}}{\sqrt{x^{4}+a^{4}}} \right)\]?

OpenStudy (anonymous):

its a^4 not a^2 :D

OpenStudy (anonymous):

yea thts it

OpenStudy (apoorvk):

so, let x^2 = a^2 sinM then 2xdx = a^2cosM.dM Use this substituion. you 'll land up with: helps?